# Demonstration of a reconfigurable ground to space quantum networking scheme

### v1

@@ Appendix A Materials and Methods
The total optical loss in the experimental setup arises from several distinct stages. At the chip interface, light was coupled into and out of the waveguide with 532 nm and 1550 nm polarization-maintaining lensed fibres (OZ Optics), incurring a coupling loss of approximately 4.5 dB. Following this, a free space bridge separated the signal and idler photons, with measured coupling efficiencies of 44 (3.57 dB loss) for idler photons and 16 (7.96 dB loss) for signal photons. The FTM unit introduced a significant insertion loss of 10 dB, due to the loss of 3 dB from the optical circulator and 7 dB attributed to the fibre Bragg grating (FBG) array. On the idler path, demultiplexing was performed using a cascaded assembly of circulators and FBGs, resulting in insertion losses of 4.86 dB, 7.73 dB, 10.48 dB, and 8.51 dB for the Alice, Bob, Charlie, and Dana channels, respectively. The 1550/780 WDM unit used for the ground configuration in Fig. 1b had an insertion loss of 0.97 and 3.02 dB at 1550 and 780 nm respectively.

### v2

@@ Appendix A Materials and Methods
Waveguide coupling
To enable precise positional adjustment, the input (OZ Optics 532 nm PM fibre) and output (1550 nm PM lensed fibre) fibres were mounted on on three-axis translation stages (MAX313D, Thorlabs), each controlled by a piezo controller (BPC203, Thorlabs). The input fibre was flat cleaved and butt coupled to the waveguide which was mounted in its oven (HC Photonics) and placed on a two-axis (vertical and horizontal) translation stage for precise positioning. The coupling process was monitored via an overhead stereo microscope (Leica M125 C) with a tunable magnification between 12.8 and 160X [43]. Coupling into and out of the waveguide introduced 4.5 dB loss. The waveguide used in the experiment was a PPLN ridge waveguide fabricated on a z-cut substrate. The chip dimensions were mm (length width thickness), with the waveguide defined on the z-facet at a depth of approximately m. The width of the channel was m and the mode field diameter were and m for the transverse y- and z-directions respectively. Following the down conversion, a free-space bridge separated the signal and idler photons, with measured coupling efficiencies of 44 (3.57 dB loss) for idler photons and 16 (7.96 dB loss) for signal photons. The reduced coupling efficiency for the latter was due to spatial mode filtering from the PM780 fibre (Thorlabs) as the nm light is slightly multi-modal in the nm fibre at the output of the waveguide.
Frequency-to-time mapping
The FTM unit consists of an array of four fibre Bragg gratings, each designed to reflect one spectral channel as indicated in Fig. 3a. More precisely, the four FBGs (O/E Land Inc.) have central wavelengths at and nm, each with a bandwidth of nm, and separated by cm of PM-780-HP fibre. As each grating is positioned at a different location along the fibre, the reflected photons traverse an additional round-trip optical path before returning, resulting in a wavelength-dependent time delay of approximately ns between adjacent channels. By mapping each frequency channel to a distinct arrival time, the FTM unit enables the satellite node to resolve multiple channels with a single detector, reducing hardware complexity while maintaining full channel discrimination. The FTM unit added 10 dB loss (3 dB from the optical circulator and 7 dB from the FBG array). For the satellite configuration, the FTM unit and the Si-APD are connected via a 5 meter 780 PM fiber (Thorlabs 780PM). Taking into account the detection efficiency, the total loss of the signal channel for the satellite configuration is dB.
Wavelength division demultiplexing
The ground users are separated by 10 meters of telecom optical fiber (Thorlabs SMF-28). Demultiplexing for the idler photons was implemented via a cascaded assembly of circulators and FBGs, resulting in losses of 4.9 dB, 7.7 dB, 10.5 dB, and 8.5 dB for Alice, Bob, Charlie, and Dana, respectively. Adding the 8 dB contribution from the waveguide loss and free-space bridge, the total idler-arm losses for the satellite configuration amount to approximately 12.9 dB, 15.7 dB, 18.5 dB, and 16.5 dB for Alice, Bob, Charlie, and Dana, respectively. For the ground configuration, the 1550/780 WDM had an insertion loss of 0.97 and 3.02 dB at 1550 and 780 nm respectively. The individual signal fiber Bragg gratings () had a loss of dB. The Si-APD had an efficiency at 780 nm. The SNSPD had an efficiency of at 1550 nm but just at 780 nm. However SNSPDs that are efficient at both wavelengths are already under development. [12, 11].